# SHaRe-SSM: An Oscillatory Spiking Neural Network for Target Variable Modeling in Long Sequences

### v1

@@ Abstract
In recent years, with the emergence of large models, there has been a significant interest in spiking neural networks (SNNs) primarily due to their energy efficiency, multiplication-free, and sparse event-based deep learning. Similarly, state space models (SSMs) in varying designs have evolved as a powerful alternative to transformers for target modeling in long sequences, thereby overcoming the quadratic dependence on sequence length of a transformer. Inspired by this progress, we here design SHaRe-SSM (Spiking Harmonic Resonate and Fire State Space Model), for target variable modeling (including both classification and regression) for very-long-range sequences. Our second-order spiking SSM, on average, performs better than transformers or first-order SSMs while circumventing multiplication operations, making it ideal for resource-constrained applications. The proposed block consumes $73\times$ less energy than second-order ANN-based SSMs for an 18k sequence, while retaining performance. To ensure learnability over the long-range sequences, we propose exploiting the stable and efficient implementation of the dynamical system using parallel scans. Moreover, for the first time, we propose a kernel-based spiking regressor using resonate and fire neurons for very long-range sequences. Our network shows superior performance on even a 50k sequence while being significantly energy-efficient. In addition, we conducted a systematic analysis of the impact of heterogeneity, dissipation, and conservation in resonate-and-fire SSMs.
@@ 1 Introduction
Spike-based deep learning is a well-established approach for efficient Artificial Intelligence, especially due to its traits of ultra-low energy consumption and sparse computation. Whether it is digital spiking neuromorphic hardware such as Loihi loihi ; ShresthaLoihi2 , TrueNorth truenorth , or analog neuromorphic hardware such as Dynapse richter2024dynap , the spike-based computation resource utilization is significantly lower than artificial neural networks.
Presently, a significant amount of these spike-based deep learning zhou2023spikformer ; 10890026 ; lee2025spikingtransformerspatialtemporalattention ; spikingssm ; stan2024learning focus on using simple integrate and fire (IF) neurons (or their leaky version (LIF)). Although IF neurons are biologically plausible, they miss key biological characteristics such as oscillatory dynamics. Such rich dynamics are observed empirically and in biophysically detailed Hodgkin-Huxley (HH) models Hodgkin1952-sd . Although the HH model provides an accurate and detailed representation of neuronal dynamics, its high computational cost makes it impractical for large-scale AI systems. But Resonate and Fire (RF) neurons izhikevich2001resonate , though computationally equivalent to IF neurons, are expressive enough to capture such rich dynamics of biological neurons. Hence, recently, there has been a significant interest in designing RF-based spiking neural networks ShresthaLoihi2 ; higuchi2024balanced ; c-silif . To the best of our knowledge, RF neurons have not been studied in a very long sequence for target variable modeling (classification as well as regression).
For sequential data processing, transformers are the de facto standard attention . But a vanilla transformer requires quadratic computational complexity with respect to the sequence length. To obviate these limitations, methods like KV caching brandon2024reducing or test-time memory updates behrouz2024titans have been proposed. However, they do not match the computational simplicity of recurrent neural networks (RNNs). Hence, there has been significant interest in designing SSMs s4 ; s5 ; mamba , including spiking neurons stan2024learning ; spikingssm ; pspikessm , which promise to get the best of both transformers and RNNs. But the existing spiking SSMs are not optimal for target variable modeling in very long sequences.
Furthermore, as detailed in (rusch2025oscillatory, ), it’s very difficult to select a state matrix for very long interactions in a first-order SSM. To overcome these constraints, they proposed LinOSS, which is a second-order state space model rusch2025oscillatory ; dlinoss . Interestingly, LinOSS attains high classification for extremely long sequence classification and regression only with a diagonal state matrix. It achieves this by leveraging stable discretizations whose spectral radius stays close to one. LinOSS is state-of-the-art in performance on extremely long-range interactions. However, it lacks key biological features, such as spike-based communication, which has also been shown to demonstrate superior energy benefits in AI applications ShresthaLoihi2 ; Imam2020-jm .
Inspired by this progress, in this work, we propose SHaRe-SSM: Spiking Harmonic Resonate and Fire State Space Model, a second-order spike-based SSM primarily targeted for extremely long-range performance. Due to spike-based computation, SHaRe-SSM is energy efficient and hence is ideal for resource-constrained applications, such as in always-on edge AI devices. In summary, our work makes the following contributions:
  • We propose and rigorously evaluate a second-order spiking SSM: SHaRe-SSM. Noteworthily, unlike some of the previous first-order SSMs pspikessm ; spikingssm ; stan2024learning , SHaRe-SSM is completely spike-based without using non-linearities such as GeLU, GLU, or GSU.
  • As part of it, for rapid training and inference, we propose a compatible parallel scan algorithm.
  • Furthermore, we observe that SHaRE-SSM outperforms all first-order SSMs on very long sequence classification. And, SHaRE-SSM is significantly energy efficient than ANN-based second-order SSMs.
  • Also, we extend SHaRe-SSM to energy-efficient long-range regression applications and propose a convolving kernel-based regression. We observe that SHaRe-SSM outperforms first-order SSMs on a 50K sequence length regression task.
  • Finally, we study and explore the impact of discretization methods and network heterogeneities on SHaRe-SSM.
2 Related Works
State Space Models:
Structured State Space Models (SSMs), including S4, Mamba, and LRU, have demonstrated state-of-the-art performance on a wide range of sequence modeling tasks, particularly in language modeling mamba and time series forecasting rusch2025oscillatory . These models leverage the mathematical structure of state space representations to model long-range dependencies efficiently and scalably s4 ; s5 ; lru . Recent work shows that neural oscillators, as universal function approximators, naturally capture long-range temporal dependencies and often outperform traditional models cornn ; unicornn ; lanthaler2023neural . They benefit from numerically stable Implicit-Explicit (IMEX) and Implicit Euler (IM) discretizations, which better handle complex dynamics than standard Explicit Euler methods. LinOSS rusch2025oscillatory builds on this oscillator-based perspective, proposing a linear implicit oscillator-based state-space model that combines the stability of implicit solvers with the expressivity of neural oscillatory dynamics. This leads to better inductive biases for learning long sequences, especially when capturing slow or multi-scale temporal phenomena. However, LinOSS is not optimal for resource-constrained applications (such as in edge AI) due to its computational complexity.
Spiking Neural Networks:
Spiking Neural Networks (SNNs) form a biologically inspired class of neural models characterized by discrete spike-based communication. The most foundational models include the IF, LIF neurons. These have been extended to include adaptation mechanisms, such as in Adaptive LIF (ALIF), and more complex dynamics as in the RF izhikevich2001resonate and Harmonic Resonate-and-Fire (HRF) hrf neurons. Many ALIF-based recurrent models have been proposed alif ; adlif ; d-cadlif , with recent works extending adaptation to RF and HRF neurons hrf ; higuchi2024balanced , which have been evaluated on only small sequence datasets. SNNs are increasingly modeled within the SSM framework for efficient computation, especially on resource-constrained hardware. (c-silif, ) reformulated LIF and RF neurons as SSMs following (adlif, ), achieving strong results on neuromorphic benchmarks. Yet, scaling to extremely long-range sequences remains a key challenge. Recent non-adaptive, first-order SSM-based spiking models such as S4D-/S4-inspired LIF neurons stan2024learning ; pspikessm ; spikingssm and RF-based variants prf perform well on mid-range tasks (up to 16k sequence length) but fall short on longer sequences. Notably, none have been evaluated on extremely long-range time series or regression benchmarks. While an FFT-based transformer wu2025spikf has been introduced for regression, its reliance on computationally heavy FFTs limits suitability for neuromorphic deployment pspikessm . (meyer2025diagonal, ) recently validated the efficacy of an S4D-based spiking SSM by implementing it on Loihi 2. Although this design was efficient, it was evaluated on reasonably simple sCIFAR-10 datasets only.
(perez-nieves_neural_2021, ; shen2024rethinking, ) showed that in LIF/IF-based networks, neural heterogeneity—variation in neuron hyperparameters can improve classification performance. In SHaRe-SSM, we aim to explore this further by focusing on HRF neurons in a second-order SSM framework, thereby investigating the combined effects of heterogeneity and discretization. Most prior studies have relied on explicit discretization schemes, with very few exploring alternatives. A notable exception is (se-adlif, ), which demonstrated improved performance using a symplectic implicit-explicit (IMEX) scheme over conventional adaptive-LIF models adlif . We here focus on enhancing the expressivity of HRF neurons by integrating them into a fully spike-based SSM framework (SHaRe-SSM).
@@ 3.1 Network Description
The Resonate-and-Fire (RF) neuron, introduced by izhikevich2001resonate , better approximates the Hodgkin-Huxley (HH) model than the Leaky Integrate-and-Fire (LIF) neuron. This is because it captures complex-valued subthreshold dynamics using a 2D linear system with a complex eigenvalue, enabling membrane resonance and frequency preference. Unlike the purely real-valued LIF neuron, RF can exhibit oscillatory behavior and selective amplification of certain input frequencies, features also observed in HH neurons. Later, second-order real-valued approximations, e.g., (hrf, ; higuchi2024balanced, ), simplify RF into a real harmonic-oscillator-like form often referred to as HRF. HRF preserves oscillatory behavior via complex eigenvalues similar to the RF neuron, but without requiring initialization in complex space; The RF model’s inherent complex dynamics are a key reason it is considered more biophysically realistic. Recently, (higuchi2024balanced, ) showed RF-based models work best without resetting the membrane potential. After collating these observations, we here integrate HRF into an SSM framework and compute neuronal dynamics linearly using parallel scans, and use the spike function as an activation function to propagate spikes.
A Harmonic-Resonate-and-Fire Neuron is defined by:
$\begin{split}u^{\prime}(t)&=-\omega^{2}v(t)-2bu(t)+x(t)\\ v^{\prime}(t)&=u(t)\\ z(t)&=\Theta(u(t)-\theta)\end{split}$ (1)
where $v(t)$ denote the hidden state, $u(t)$ as the time derivative of $v(t)$, $x(t)$ is the input spikes to the neuron, $\omega\in\mathbb{R}$ is the frequency parameter and $b\in\mathbb{R}$ is the damping coefficient. Spiking Function: $\Theta\in\{0,1\}$ i.e. spikes are emitted if $u\geq\theta$.
Figure 1: Description of SHaRe-SSM model. SHaRe-SSM consists of a SHaRe-SSM neuron representing multiple states (Eq. 1) followed by a Linear layer repeated for N blocks. Spikes are encoded with a Spike Encoder and decoded using a Spike Decoder.
This HRF formulation is limited to a single neuron. We propose formulating HRF in an SSM framework (SHaRe-SSM). This approach helps in capturing better dynamics. The input spikes are multiplied by a weight matrix, and we propagate spikes from the hidden state $v(t)$. Moreover, removing the damping parameter in such second-order approximations enhances the neuron’s ability to capture long-range temporal dependencies by preserving longer oscillations rusch2025oscillatory . We define the SHaRe-SSM model by:
$\begin{split}u^{\prime}(t)&=-\Omega v(t)+Bx(t)\\ v^{\prime}(t)&=u(t)\\ z(t)&=\Theta(v(t)-\theta)\\ y(t)&=Cz(t)+Dx(t)\end{split}$ (2)
@@ A Harmonic-Resonate-and-Fire Neuron is defined by:
where $u(t),v(t)\in\mathbb{R}^{p}$ denote the hidden states, $y(t)\in\mathbb{R}^{h}$ the output, and $x(t)\in\mathbb{R}^{h}$ the input spike signal. The system is defined by weights $\Omega\in\mathbb{R}^{p\times p}$ which is diagonal, $B\in\mathbb{R}^{p\times h}$, $C\in\mathbb{R}^{h\times p}$, $D\in\mathbb{R}^{h}$, and an output learnable spiking parameter $\theta\in\mathbb{R}^{p}$ such that spikes are emitted if $v\geq\theta$. These thresholds ($\theta$) are learned using a step-double Gaussian surrogate gradient neftci2019surrogate ; higuchi2024balanced . Since higuchi2024balanced showed that HRF works well with no reset, we here employed the same strategy. This avoids the need for sequential processing, and the model can be implemented across time as an activation function. Figure 1 describes the SHaRe-SSM design. Each SHaRe-SSM block is comprised of SHaRe-SSM neuron, followed by a linear layer, batch normalization, and an IF neuron. We mix spikes from the previous layer and introduce dropout layers to avoid learnable thresholds from saturating early. SHaRe-SSM uses only binary spike-based communication between blocks.
@@ 3.2 Encoder and Decoder
For SHaRe-SSM, we design an encoder for encoding the input signal into spike trains (see Figure 1). Herein, the inputs every timestep are passed through a learnable linear layer. The output of this layer is passed through a batch normalization layer. Finally, an IF neuron generates spikes into the SHaRe-SSM block. Notably, this is a data-dependent trainable encoder (without requiring us to specify whether to utilize rate coding or any other encoding variants). Moreover, it utilizes neuronal heterogeneity and is instantaneous.
The decoder projects the spikes back using a linear Layer with the output dimension equal to the number of classes. For regression, we also propose convolving the output with a learnable filter.
3.3 Discretization methods
For understanding the impact of discretization on an HRF neuron (Eq. 1), we conduct a single neuron simulation (Figure 2) by injecting a unit spike input to an HRF neuron with zero phase. For the continuous case, if the damping parameter ’b’ is constrained to be non-negative, we must get stable dynamics. If set to zero, we should observe sustained oscillation. From Figure 2, we can see that for Euler Forward (Explicit), our model is diverging over time, making it unstable. IMEX discretization is overall stable and hence preserves energy. IM, although dissipative in energy, is still stable. We briefly describe IM and IMEX generalizability for long sequences in detail below.
Figure 2: Post-spike membrane potential behavior for a Harmonic RF neuron for different discretizations with no damping. The dotted red line denotes the inter-spike interval.
IMEX is promising, as (rusch2025oscillatory, ) showed that removing damping leads to a Hamiltonian system arnold1989mathematical , where energy conservation is ensured under symplectic discretizations. In addition, we examine the non-conservative but stable IM scheme, which performs well empirically. We adopt a second-order system with position-like state $u_{n}$ and velocity-like state $v_{n}$. The general form of discrete updates is denoted by:
@@ 3.3 Discretization methods
where the choice of $v_{\star}$ distinguishes the discretization schemes: $v_{\star}=v_{n}$ for IM (implicit), and $v_{\star}=v_{n-1}$ for IMEX (implicit-explicit).
For the IM scheme, we define rusch2025oscillatory :
$M^{\text{IM}}$ $=\begin{pmatrix}S&-S\Delta t\Omega\\ S\Delta t&S\end{pmatrix},\quad F_{n}^{\text{IM}}=\begin{pmatrix}S\Delta tBx_{n}\\ S\Delta t^{2}Bx_{n}\end{pmatrix}$,
with $S=(I+\Delta t^{2}\Omega)^{-1}$, acting as a Schur complement that suppresses high-frequency components and ensures eigenvalues of $M^{\text{IM}}$ remain bounded within the unit circle.
And, for the IMEX scheme, we use rusch2025oscillatory :
$M^{\text{IMEX}}$ $=\begin{pmatrix}I&-\Delta t\Omega\\ \Delta tI&I-\Delta t^{2}\Omega\end{pmatrix},\quad F_{n}^{\text{IMEX}}=\begin{pmatrix}\Delta tBx_{n}\\ \Delta t^{2}Bx_{n}\end{pmatrix}$.
Unlike IM, the eigenvalues of $M^{\text{IMEX}}$ lie near the unit circle, preserving oscillatory energy over long horizons. This makes IMEX especially suitable for modeling long sequences, where maintaining temporal structure is crucial.
3.4 Parallel scan in SHaRe-SSM
Parallel scans kogge1973parallel ; blelloch1990prefix exploit associativity to reduce the time complexity of recurrent computations from $\mathcal{O}(N)$ to $\mathcal{O}(\log N)$. Originally developed to accelerate RNNs, they have recently been adapted for state-space models s5 , powering modern architectures like Linear Recurrent Units (LRUs) lru and Mamba mamba . Leveraging this structure allows for richer exploration of neuronal state trajectories and improves training efficiency. We apply parallel scans to the linear computations, followed by spike function activations.
An associative binary operation similar to rusch2025oscillatory defined as:
$(a_{1},a_{2})\bullet(b_{1},b_{2})=(b_{1}\cdot a_{1},b_{1}\cdot a_{2}+b_{2}),$
where $\cdot$ represents matrix-matrix and matrix-vector multiplications. Applying a parallel scan to the input sequence $\left[(M,F_{1}),(M,F_{2}),\ldots\right]$ efficiently computes the solution to the $s_{n}=Ms_{n-1}+F_{n}$. The scan output stores $x_{n}$ in the second element of each tuple. This is made efficient by exploiting structured matrices (e.g., diagonal block $2\times 2$ matrices in $M_{IM}$ and $M_{IMEX}$), where each product can be computed in linear time with respect to the hidden dimension. We use this approach to leverage IMEX and IM discretization schemes on an HRF neuron.
Algorithm 1 describes our SHaRE-SSM implementation.
@@ 3.4 Parallel scan in SHaRe-SSM
  $y^{n}\leftarrow drop(\Theta(y^{n}-\theta_{D}^{n}))$
@@
  $y^{n}\leftarrow\textit{BN}(y^{n}))$
  $y^{n}\leftarrow drop(\Theta(y^{n}-\theta^{n})))$
  $x^{n}\leftarrow cat(x^{n-1},y^{n})$

### v2

# Learning from Time Series for Health

@@ Abstract
Spiking neural networks have garnered increasing attention due to their energy efficiency, multiplication-free computation, and sparse event-based processing. In parallel, state space models have emerged as scalable alternatives to transformers for long-range sequence modelling by avoiding quadratic dependence on sequence length. We propose SHaRe-SSM (Spiking Harmonic Resonate-and-Fire State Space Model), a second-order spiking SSM for classification and regression on ultra-long sequences. SHaRe-SSM outperforms transformers and first-order SSMs on average while eliminating matrix multiplications, making it highly suitable for resource-constrained applications. To ensure fast computation over tens of thousands of time steps, we leverage a parallel scan formulation of the underlying dynamical system. Furthermore, we introduce a kernel-based spiking regressor, which enables the accurate modelling of dependencies in sequences of up to 50k steps. Our results demonstrate that SHaRe-SSM achieves superior long-range modelling capability with energy efficiency ($52.1\times$ less than ANN-based second-order SSM), positioning it as a strong candidate for resource-constrained devices such as wearables.
@@ 1 Introduction
Spike-based deep learning has established itself as an ultra-low-power consumption and sparse computing paradigm for efficient AI in recent years. Spike-based neuromorphic hardware such as Loihi [loihi, ShresthaLoihi2], TrueNorth [truenorth], and Dynapse [richter2024dynap] utilises far lower resources than conventional ANN-based designs. Apparently, most spike-based models [zhou2023spikformer, 10890026, lee2025spikingtransformerspatialtemporalattention, spikingssm, stan2024learning] rely on integrate-and-fire (IF) or leaky IF neurons, which miss key biological traits like oscillations. While the biophysically detailed Hodgkin-Huxley model captures these dynamics, it is computationally prohibitive. Driven by these observations, Resonate-and-fire (RF) neurons [izhikevich2001resonate], computationally as light as IF but more expressive, have gained recent attention [ShresthaLoihi2, higuchi2024balanced, c-silif]. However, RF neurons remain underexplored for very long sequence modelling.
For sequential tasks, transformers are de facto standards [attention], but they suffer from quadratic dependence on sequence length. Alternatives, such as KV caching [brandon2024reducing] and memory updates [behrouz2024titans], reduce overhead but lack the simplicity of RNNs. State space models (SSMs) [s4, s5, mamba] and their spiking variants [stan2024learning, spikingssm, pspikessm] bridge this gap. Yet current spiking SSMs struggle with very long sequences.
LinOSS [rusch2025oscillatory, dlinoss], a second-order SSM, addresses this by using stable discretisations with diagonal state matrices, achieving state-of-the-art results on long-range tasks. However, it lacks spike-based communication, which is crucial for energy efficiency [ShresthaLoihi2, Imam2020-jm], such as demanded by a battery-driven wearable. Hence, in this work, we introduce SHaRe-SSM, a second-order spiking SSM designed for extremely long-range tasks and energy-efficient edge AI. Our contributions are: (1) A fully spike-based second-order SSM, without ANN nonlinearities (GeLU, GLU, or GSU). (2) A compatible parallel scan algorithm for fast training and inference. Superior accuracy over first-order SSMs on long-sequence classification, with higher efficiency than ANN-based second-order SSMs. (3) An extension to regression via a convolving kernel, outperforming first-order SSMs on 50K-length tasks.
@@ 2.1 Network Description
The Resonate-and-Fire (RF) neuron [izhikevich2001resonate] provides a closer approximation to the Hodgkin-Huxley (HH) model than the Leaky Integrate-and-Fire (LIF) neuron by capturing subthreshold resonance through a 2D linear system with complex eigenvalues. This enables oscillatory dynamics and frequency selectivity, traits absent in LIF but observed in HH neurons. To simplify implementation, second-order real-valued variants such as the Harmonic RF (HRF) [hrf, higuchi2024balanced] reformulate RF as a real harmonic oscillator while preserving oscillatory behaviour without requiring complex initialisation. RF’s complex dynamics make it more biophysically realistic, and recent work [higuchi2024balanced] shows that RF-based models perform best without membrane resets. Building on these insights, we integrate HRF into LinOSS’s SSM framework [rusch2025oscillatory], compute dynamics linearly via parallel scans, and use a spike function as the activation to propagate spikes. This original HRF formulation is limited to a single neuron. We propose formulating HRF in an SSM framework (SHaRe-SSM). This approach helps in capturing better dynamics. The input spikes are multiplied by a weight matrix, and we propagate spikes from the hidden state $v(t)$. Moreover, removing the damping parameter in such second-order approximations enhances the neuron’s ability to capture long-range temporal dependencies by preserving longer oscillations [rusch2025oscillatory]. We define the SHaRe-SSM model by:
[…]
@@
where $u(t),v(t)\in\mathbb{R}^{p}$ denote the hidden states, $y(t)\in\mathbb{R}^{h}$ the output, and $x(t)\in\mathbb{R}^{h}$ the input spike signal. The system is defined by weights $\Omega\in\mathbb{R}^{p\times p}$ which is diagonal, $B\in\mathbb{R}^{p\times h}$, $C\in\mathbb{R}^{h\times p}$, $D\in\mathbb{R}^{h}$, and an output learnable spiking parameter $\theta_{C}\in\mathbb{R}^{p}$ such that spikes are emitted if $v\geq\theta_{C}$. These thresholds ($\theta$) are learned using a step-double Gaussian surrogate gradient [neftci2019surrogate, higuchi2024balanced]. Since [higuchi2024balanced] demonstrated that HRF performs well without a reset, we adopted the same strategy here. This eliminates the need for sequential processing, allowing the model to be implemented across time as an activation function. Each SHaRe-SSM block is comprised of a SHaRe-SSM neuron, followed by a linear layer, and a spike function (IF neuron with no reset).
@@ 2.2 Encoder and Decoder
For SHaRe-SSM, we design an encoder for encoding the input signal into spike trains (see Algorithm 1). Herein, the inputs every timestep are passed through a learnable linear layer. Finally, an IF neuron with no reset, i.e., a spike function, generates spikes into the SHaRe-SSM block. Notably, this is a data-dependent trainable encoder (without requiring us to specify whether to utilise rate coding or any other encoding variants). Moreover, it utilises neuronal heterogeneity and is instantaneous. The decoder projects the spikes back using a linear Layer with the output dimension equal to the number of classes. For regression, we also propose convolving the output with a learnable filter.
2.3 Discretisation methods
Euler Forward (explicit) discretisation causes divergence over time, making the model unstable. In contrast, IMEX discretisation remains stable and preserves energy, while the IM scheme, though dissipative, is also stable. Below, we analyse the generalizability of IM and IMEX for long sequences. IMEX is particularly promising for regression: as shown in rusch2025oscillatory, removing damping yields a Hamiltonian system [arnold1989mathematical], where energy conservation is guaranteed under symplectic discretisations. We also evaluate the non-conservative but empirically stable IM scheme. Our formulation adopts a second-order system with position-like state $u_{n}$ and velocity-like state $v_{n}$. The general form of discrete updates is denoted by:
@@ 2.3 Discretisation methods
where the choice of $v_{\star}$ distinguishes the discretisation schemes: $v_{\star}=v_{n}$ for IM (implicit), and $v_{\star}=v_{n-1}$ for IMEX (implicit-explicit).
Unlike IM, the eigenvalues of $M^{\text{IMEX}}$ lie near the unit circle, preserving oscillatory energy over long horizons. This makes IMEX especially suitable for modelling long sequences, where maintaining temporal structure is crucial.
@@
  $y^{n}\leftarrow\Theta(y^{n}-\theta_{D}^{n})$
@@
  $y^{n}\leftarrow\Theta(y^{n}-\theta^{n})$
  $x^{n}\leftarrow y^{n}+x^{n-1}$ {Spike mixing}